Why does the mass of an object increase when its speed approaches that of light?

Grade:9

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
The relativistic mass of an object is given by
m=\frac{m_{0}}{\sqrt{1-\frac{v^{2}}{c^{2}}}}
Therefore, as the velocity increases, and approaches the speed of light, the mass keeps on increasing and approaches infinity.
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty